# 3DRot: 3D Rotation Augmentation for RGB-Based 3D Tasks

### v2

@@ Abstract
RGB-based 3D tasks, e.g., 3D detection, depth estimation, 3D keypoint estimation, still suffer from scarce, expensive annotations and a thin augmentation toolbox, since most image transforms, including resize and rotation, disrupt geometric consistency. In this paper, we introduce 3DRot, a plug-and-play augmentation that rotates and mirrors images about the camera’s optical center while synchronously updating RGB images, camera intrinsics, object poses, and 3D annotations to preserve projective geometry-achieving geometry-consistent rotations and reflections without relying on any scene depth. We validate 3DRot with a classical 3D task, monocular 3D detection. On SUN RGB-D dataset, 3DRot raises $IoU_{3D}$ from 43.21 to 44.51, cuts rotation error (ROT) from 22.91 to 20.93, and boosts $mAP_{0.5}$ from 35.70 to 38.11. As a comparison, Cube R-CNN adds 3 other datasets together with SUN RGB-D for monocular 3D estimation, with a similar mechanism and test dataset, increases $IoU_{3D}$ from 36.2 to 37.8, boosts $mAP_{0.5}$ from 34.7 to 35.4. Because it operates purely through camera-space transforms, 3DRot is readily transferable to other 3D tasks.
11footnotetext: This work was done during an internship at the International Digital Economy Academy (IDEA).33footnotetext: Corresponding author.
Introduction
@@
RGB-based 3D perception has become a cornerstone of robotics, autonomous driving, and augmented-reality workflows. However, curating large-scale 3D datasets is far more expensive and time-consuming than labeling 2D images: annotators must specify metric object poses and sizes, often with laser scans or multi-view capture, and reports show that 3D boxes can cost several-fold more per instance than 2D boxes. Data scarcity consequently bottlenecks generalization and fuels the search for synthetic diversity (Brazil et al. 2023). Focusing on non-generative, non-insertion, plug-and-play augmentations, today’s RGB-based 3D detection pipelines still rely on a narrow menu-chiefly random scaling and cropping (with intrinsic updates), horizontal flips (and only rarely vertical flips), and global color jitter-because any transform must preserve geometric consistency so that augmented objects remain physically plausible in the scene (Brazil and Liu 2019; Li et al. 2022; Zou et al. 2021; Chen et al. 2020; Li, Jia, and Shi 2024). Indeed, naively pasting objects at arbitrary depths or orientations injects implausible scale-depth cues and can hurt detector accuracy (Ge et al. 2023; Li, Jia, and Shi 2024; Parihar et al. 2025).
(1)Geometry-Consistent Camera/Scene Transforms. These methods perturb the camera pose or the ground plane and re-project all annotations, thereby preserving the exact 2D–3D correspondence. Typical operations include scaling, cropping, ground–plane-constrained rotations, and flipping (Lian et al. 2022; Engilberge et al. 2023). In multi-sensor fusion and multi-view settings, flipping is a common choice as long as the transformation is applied consistently across modalities/views (Wang et al. 2021; Pan et al. 2025). The advantages are simplicity, no rendering pipeline, and plug-and-play integration; however, the operation set remains narrow, and rotation augmentation is restricted to in-plane/coplanar assumptions and thus cannot be directly used when views or data are not approximately coplanar. (2)Physically Plausible Instance/Object Insertion. These approaches cut object instances from real scenes or CAD repositories, paste them in 3D or BEV, and synchronize the RGB projections while explicitly handling occlusion and scale (Wang et al. 2021; Ge et al. 2023; Parihar et al. 2025). The main drawback is the heavy overhead of scene reconstruction and photometric harmonization (e.g., collision checks, lighting/shadow consistency), which hampers scalability across domains. (3)Synthetic / Generative and Weak-Supervision Pipelines. These pipelines first reconstruct the 3D scene (or novel viewpoints), then edit objects or cameras in the 3D world and re-render to produce training images (Wang et al. 2024; Cheng et al. 2025; Chang et al. 2024). Geometric consistency is guaranteed by the rendering process and the resulting realism can be photo-realistic; nevertheless, reconstruction and rendering are computationally expensive, and iteration cycles are slow.
@@
Recent progress has gravitated toward physically realistic instance insertion and generative/weakly supervised pipelines rather than toward truly plug-and-play primitives. Object insertion methods synthesize diversity by pasting CAD or real instances with explicit occlusion/scale handling (Ge et al. 2023; Parihar et al. 2025), while generative routes reconstruct scenes or viewpoints and then re-render (Wang et al. 2024; Cheng et al. 2025; Chang et al. 2024). Consequently, the most fundamental RGB-based setting, monocular RGB, still relies primarily on horizontal flips and mild color jitter (Brazil and Liu 2019; Li et al. 2022; Zou et al. 2021; Chen et al. 2020; Li, Jia, and Shi 2024), while in-plane rotation-an indispensable and highly effective augmentation in 2D recognition-remains largely untapped in 3D settings. However, naive rotation or pasting of objects in 3D space can destroy tight depth-scale coupling and violate geometric consistency, leading to a degraded detector accuracy (Ge et al. 2023; Li, Jia, and Shi 2024; Parihar et al. 2025). Even the ubiquitous horizontal flip still lacks a principled account of when and why it helps. These observations motivate us to devise a plug-and-play, geometry-aware augmentation module that enriches object-pose diversity without information loss, alleviates annotation scarcity, and lays the groundwork for more sophisticated, geometry-driven augmentations in future work.
@@
In this paper, we propose Optical Center 3D Rotation Augmentation (3DRot), a plug-and-play module that enriches RGB-based 3D training data by rotating and mirroring scenes about the camera’s optical center while rigorously preserving projective geometry as shown in Fig. 1. For each desired in-plane rotation or reflection, 3DRot derives an exact homography from the camera projection matrix, enabling one-to-one pixel remapping without depth hallucination; the same closed-form transformation keeps the RGB image, 3D data, and camera intrinsics mutually consistent, delivering depth-free, geometry-faithful augmentation that drops seamlessly into existing 3D tasks.
@@
As a fundamental RGB-based 3D task, monocular 3D detection serves as our representative testbed to demonstrate the effectiveness of 3DRot. We validate 3DRot by inserting it into a frozen DINO-X (Ren et al. 2024; Liu et al. 2024) + Cube R-CNN head (Brazil et al. 2023) with matched schedules. On SUN RGB-D 10-category split (SUN10) (Song, Lichtenberg, and Xiao 2015), it improves $IoU_{3D}$ from 43.21 to 44.51, lowers ROT from 22.91 to 20.93, and increases $mAP_{0.5}$ from 35.70 to 38.11; smaller but consistent gains appear on SUN10. Ablations confirm geometry-consistent rotations and chirality-safe flips as the main drivers, with principal-point realignment providing a minor additional benefit.
Related Works
@@
RGB-based Rotation Augmentation
@@
Early attempts to regularize RGB-based networks with viewpoint changes started from settings in which the depth channel is readily available, so the augmentation can be expressed directly in camera space. Keypoint-based Monocular 3D Object Detector (Kim et al. 2020) injects small perturbations into the robot‐mounted camera’s roll, pitch and yaw and then re-project SUN RGB-D boxes to match the new pose. While this improves robustness to head motions, it still relies on full RGB-D frames. Two-level Data Augmentation (Engilberge et al. 2023) pushes consistency a step further: their two-level scheme first applies an in-plane homography to each calibrated view and then corrects the per-view projection matrix so that all images still meet at a common ground plane. The operation is, however, intrinsically coplanar; any 3D structure that rises above or falls below the reference plane (e.g. shelves or hanging objects) accumulates geometric error after transformation. Most recently, GroundMix (Meier et al. 2024) shows that simple in-plane rotations around the optical axis suffice for monocular car- and drone-view detectors; they adjust 3D yaw accordingly and keep the remaining pose angles fixed. This assumption holds only when the camera truly spins about its z-axis; moreover, it offers no rigorous proof of geometric consistency, merely demonstrating visual plausibility.
RGB-based Flipping Augmentation
@@
Random horizontal flipping is the most widely used geometry-aware transform in RGB-based tasks because it is label‐preserving for the image plane and inexpensive to implement. The pioneering (Brazil and Liu 2019) pipeline reflects each 3D box corner on the image, then back-projects the flipped vertices with the original depth to obtain an updated 3D cuboid, leaving all other pose parameters. Although effective for heading predictions, this strategy cannot propagate sign changes to the off-diagonal entries of a full $3{\times}3$ rotation matrix and therefore assumes that only the yaw angle is modeled. Several follow-up methods (Chen et al. 2020; Zou et al. 2021; Li et al. 2022; Li, Jia, and Shi 2024; Wang et al. 2021; Pan et al. 2025) inherit the same augmentation. Cube R-CNN (Brazil et al. 2023) introduces a more principled update: it composes two mirror matrices, one for the image warp and one for the camera coordinate frame, so that the predicted $R\!\in\!SO(3)$ remains valid after a flip. This design finally enables horizontal mirroring for full‐angle pose regression, but the derivation presumes that the principal point coincides with the image center; when intrinsics are asymmetric or cropped, the 2D-3D alignment still drifts.
Other Augmentations
@@
Beyond rotations and flips, several geometry–aware strategies have been explored to enrich RGB-based 3D data. Geometric Consistency Augmentation (Lian et al. 2022) introduces four camera-centric manipulations-random scaling, random cropping, camera translation, and instance‐level copy-paste-each designed to preserve depth cues after image warping. While these augmentations improve robustness, random cropping can discard large portions of background context, and the copy-paste pipeline requires dense depth maps plus heuristic rules to blend pasted objects seamlessly. 3D Copy-Paste (Ge et al. 2023) makes this idea more rigorous by sampling physically plausible insertions from an external CAD library: candidate objects are resized, posed, and lit to avoid collisions and match the scene illumination, yielding more realistic composites. Nevertheless, its rendering loop and collision search add notable pre-processing cost, and downstream accuracy still depends on perfect plane reconstruction and light estimation. Moreover, random cropping risks discarding the background context that conveys important geometric cues, while instance-level copy paste-despite its popularity-relies on intricate manual placement heuristics and still falls short of producing truly photorealistic composites (Li, Jia, and Shi 2024).
Method
@@
3DRot applies an arbitrary camera‑centric rotation (pitch, roll, yaw; optional mirroring) about the optical center and warps the image using the corresponding pure‑rotation homography. This keeps updates to the camera frame and screen space consistent and preserves the projection geometry. We first outline the framework and notation, then derive the rotation model and label‑synchronization rules. Intuitively, 3DRot rotates each pixel’s viewing ray about the optical center, preserving all geometric correspondences without introducing any new quantities that would need to be re-estimated.
@@
Coordinate System & 3D BBOX Transformation
To propagate a virtual camera rotation to all scene geometry, we must first clarify how 3D points migrate between the world frame, the original camera frame, and the rotated camera frame. Accordingly, this subsection revisits the world-to-camera transform, derives the point-wise update induced by a camera-centric rotation, and then lifts the result to the cuboid pose that will be used throughout the remainder of the method.
A point expressed in the world frame is mapped to the camera frame through the following transform:
$T_{c}=R\,T_{w}+t$ (1)
where $T_{c}\in\mathbf{R}^{3\times 1}$ represents a 3D point location in camera’s coordinate system, $T_{w}\in\mathbf{R}^{3\times 1}$ represents a 3D point location in world’s coordinate system, $R\in\mathbf{R}^{3\times 3}$ and $t\in\mathbf{R}^{3\times 1}$ represent camera extrinsics. After rotating the camera, the new camera’s coordinate system should be:
$T_{rot-c}=R_{c}\,T_{c}$ (2)
Combined (1) and (2):
$T_{rot-c}=R_{c}\,R\,T_{w}+R_{c}\,t$ (3)
(3) indicates that the camera extrinsics should be updated if rotation augmentation is applied.
The previous derivation shows how any single 3‐D point is transferred from the original camera frame to the rotated one. We now extend this result to an entire axis‐aligned cuboid, whose pose is parameterized by its translation, sizes, and rotation matrix. Let
$T_{bb}=R_{bb}\,S_{bb}\;+\;t_{bb}$ (4)
denote the $3\times 8$ matrix of the eight cuboid corners in the original camera coordinates, where the column matrix $S_{bb}\in R^{3\times 8}$ is formed by concatenating the vectorized offsets $\bigl{(}\pm\frac{W}{2},\,\pm\frac{H}{2},\,\pm\frac{L}{2}\bigr{)}^{\!\top}$.
Applying the camera‐centric rotation (2) to every corner gives
$T_{bb}^{new}=R_{c}\bigl{(}R_{bb}\,S_{bb}+t_{bb}\bigr{)}=\underbrace{(R_{c}R_{bb})}_{R_{new}}S_{bb}+\underbrace{R_{c}t_{bb}}_{t_{new}},$
hence the updated pose parameters are
$R_{new}=R_{c}\,R_{bb},\qquad S_{new}=S_{bb},\qquad t_{new}=R_{c}\,t_{bb}.$
That is, the cuboid’s orientation and translation are left-multiplied by the camera rotation $R_{c}$, whereas its side lengths remain unchanged.
Projective Transformation
Having established how the 3D coordinates of each scene point are updated under a camera-centric rotation, we now turn to their image-plane projections. Specifically, we must determine how every pixel location transforms so that the rotated view remains geometrically consistent with the new camera frame. To this end, we (i) recap the pinhole-projection equation, (ii) derive the homography that arises when the camera rotates about its optical center, and (iii) show how this homography simplifies to a pure-rotation form that is valid for any 3D scene, without the usual planarity assumption or depth information. These steps yield a closed-form mapping $H=K^{\prime}R_{c}K^{-1}$ which we will use to warp the RGB image, masks, and NOCS maps (Krishnan et al. 2024) in the remainder of the pipeline.

### v3

# 3DRot: Rediscovering the Missing Primitive for RGB-Based 3D Augmentation

@@ Abstract
RGB-based 3D tasks, e.g., 3D detection, depth estimation, 3D keypoint estimation, still suffer from scarce, expensive annotations and a thin augmentation toolbox, since many image transforms, including rotations and warps, disrupt geometric consistency. While horizontal flipping and color jitter are standard, rigorous 3D rotation augmentation has surprisingly remained absent from RGB-based pipelines, largely due to the misconception that it requires scene depth or scene reconstruction. In this paper, we introduce 3DRot, a plug-and-play augmentation that rotates and mirrors images about the camera’s optical center while synchronously updating RGB images, camera intrinsics, object poses, and 3D annotations to preserve projective geometry, achieving geometry-consistent rotations and reflections without relying on any scene depth. We first validate 3DRot on a classical RGB-based 3D task, monocular 3D detection. On SUN RGB-D, inserting 3DRot into a frozen DINO-X + Cube R-CNN pipeline raises $IoU_{3D}$ from 43.21 to 44.51, cuts rotation error (ROT) from 22.91 to 20.93, and boosts $mAP_{0.5}$ from 35.70 to 38.11; smaller but consistent gains appear on a cross-domain IN10 split. Beyond monocular detection, adding 3DRot on top of the standard BTS augmentation schedule further improves NYU Depth v2 from 0.1783 to 0.1685 in abs-rel (and 0.7472 to 0.7548 in $\delta<1.25$), and reduces cross-dataset error on SUN RGB-D. On KITTI, applying the same camera-centric rotations in MVX-Net (LiDAR+RGB) raises moderate 3D AP from about 63.85 to 65.16 while remaining compatible with standard 3D augmentations.
3D perception, data augmentation, projective geometry
{syang214@jh.edu, 2300432059@email.szu.edu.cn, jiangxiaoke@idea.edu.cn, leizhang@idea.edu.cn}
1 Introduction
RGB-based 3D perception has become a cornerstone of robotics, autonomous driving, and augmented-reality workflows. However, curating large-scale 3D datasets is far more expensive and time-consuming than labeling 2D images: annotators must specify metric object poses and sizes, often with laser scans or multi-view capture, and reports show that 3D boxes can cost several-fold more per instance than 2D boxes. Data scarcity consequently bottlenecks generalization and fuels the search for synthetic diversity (Brazil et al., 2023). Focusing on non-generative, non-insertion, plug-and-play augmentations, today’s RGB-based 3D detection pipelines still rely on a narrow menu-chiefly random scaling and cropping (with intrinsic updates), horizontal flips (and only rarely vertical flips), and global color jitter-because any transform must preserve geometric consistency so that augmented objects remain physically plausible in the scene (Brazil and Liu, 2019; Chen et al., 2020; Zou et al., 2021; Li et al., 2022, 2024; Pu et al., 2025; Piccinelli et al., 2025). Indeed, naively pasting objects at arbitrary depths or orientations injects implausible scale-depth cues and can hurt detector accuracy (Ge et al., 2023; Li et al., 2024; Parihar et al., 2025).
(1) Geometry-Consistent Camera/Scene Transforms. These methods perturb the camera pose or the ground plane and re-project all annotations, thereby preserving the exact 2D–3D correspondence. Typical operations include scaling, cropping, ground–plane-constrained rotations, and flipping (Engilberge et al., 2023; Lian et al., 2022). In multi-sensor fusion and multi-view settings, flipping is a common choice as long as the transformation is applied consistently across modalities/views (Wang et al., 2021; Pan et al., 2025a). The advantages are simplicity, no rendering pipeline, and plug-and-play integration; however, the operation set remains narrow, and rotation augmentation is restricted to in-plane/coplanar assumptions and thus cannot be directly used when views or data are not approximately coplanar. (2) Physically Plausible Instance/Object Insertion. These approaches cut object instances from real scenes or CAD repositories, paste them in 3D or BEV, and synchronize the RGB projections while explicitly handling occlusion and scale (Ge et al., 2023; Parihar et al., 2025; Wang et al., 2021). The main drawback is the heavy overhead of scene reconstruction and photometric harmonization (e.g., collision checks, lighting/shadow consistency), which hampers scalability across domains. (3) Synthetic / Generative and Weak-Supervision Pipelines. These pipelines first reconstruct the 3D scene (or novel viewpoints), then edit objects or cameras in the 3D world and re-render to produce training images (Wang et al., 2024; Cheng et al., 2025; Chang et al., 2024). Geometric consistency is guaranteed by the rendering process and the resulting realism can be photo-realistic; nevertheless, reconstruction and rendering are computationally expensive, and iteration cycles are slow.
@@ 1 Introduction
Figure 2: Depth maps. The block is a $2\times 2$ grid shown in row-major order with camera rotations about the optical center: $(0^{\circ},0^{\circ},0^{\circ})$, yaw $40^{\circ}$, pitch $20^{\circ}$, and roll $20^{\circ}$. In all cases 3DRot applies the same pure-rotation homography to the RGB image and updates labels/intrinsics accordingly, so the depth remains 2D–3D consistent while the image footprint changes. See Supplementary Video for an animated demo.
@@
Recent progress has gravitated toward physically realistic instance insertion and generative/weakly supervised pipelines rather than toward truly plug-and-play primitives. Object insertion methods synthesize diversity by pasting CAD or real instances with explicit occlusion/scale handling (Ge et al., 2023; Parihar et al., 2025), while generative routes reconstruct scenes or viewpoints and then re-render (Wang et al., 2024; Cheng et al., 2025; Chang et al., 2024). Consequently, the most fundamental RGB-based setting, monocular RGB, still relies primarily on horizontal flips and mild color jitter (Brazil and Liu, 2019; Chen et al., 2020; Zou et al., 2021; Li et al., 2022, 2024; Pu et al., 2025; Piccinelli et al., 2025), while in-plane rotation-an indispensable and highly effective augmentation in 2D recognition-remains largely untapped in 3D settings. Even the ubiquitous horizontal flip still lacks a principled account of when and why it helps. These observations motivate us to devise a plug-and-play, geometry-aware augmentation module that enriches object-pose diversity without information loss, alleviates annotation scarcity, and lays the groundwork for more sophisticated, geometry-driven augmentations in future work.
To the best of our knowledge, despite its conceptual simplicity, a rigorous optical-center rotation primitive remains absent from the literature. It is missing from standard augmentation toolboxes (Contributors, 2020) and recent RGB-based pipelines(Brazil et al., 2023; Li et al., 2024; Pu et al., 2025; Piccinelli et al., 2025), and even methods focused on geometric consistency (Lian et al., 2022) do not categorize or employ it. We hypothesize that this gap exists because the crucial insight—that a depth-free homography can perfectly synchronize intrinsics, chirality, and labels without the limiting assumptions of standard heuristics—is easily overlooked.
@@
In this paper, we propose Optical Center 3D Rotation Augmentation (3DRot), a plug-and-play module that enriches RGB-based 3D training data by rotating and mirroring scenes about the camera’s optical center while rigorously preserving projective geometry as shown in Fig. 1. For each desired in-plane rotation or reflection, 3DRot derives an exact homography from the camera projection matrix, enabling pixel remapping with no need to simulate occlusions, shadows, or lighting; the same closed-form transformation keeps the RGB image, 3D data, and camera intrinsics mutually consistent, delivering depth-free, geometry-faithful augmentation that drops seamlessly into existing 3D tasks. Moreover, multi-modal signals (e.g., LiDAR point clouds and depth-sensor maps) can be updated in lockstep with the RGB using the same camera-centric rotation/reflection, addressing the long-standing issue of cross-modal augmentation asynchrony in 3D detection and related tasks (Zhang et al., 2020; Wang et al., 2021; Li et al., 2023; Pan et al., 2025b) ; see Fig. 2 and  4 for depth-map and NOCS (Krishnan et al., 2024) examples. The formulation is especially pertinent for platforms with rapidly changing orientations—e.g., UAVs, aerospace imaging payloads, and dynamic robots—where robustness to roll/pitch/yaw is crucial. We validate 3DRot on monocular 3D cuboid detection, monocular depth estimation, and LiDAR+RGB detection, demonstrating that it is task- and modality-agnostic.
@@
As a fundamental RGB-based 3D task, monocular 3D detection serves as our main testbed. We insert 3DRot into a frozen DINO-X (Ren et al., 2024; Liu et al., 2024) + Cube R-CNN head (Brazil et al., 2023) with matched schedules. On the SUN RGB-D 10-category split (SUN10) (Song et al., 2015), 3DRot improves $IoU_{3D}$ from 43.21 to 44.51, lowers ROT from 22.91 to 20.93, and increases $mAP_{0.5}$ from 35.70 to 38.11; smaller but consistent gains appear on a cross-domain IN10 split. Ablations confirm geometry-consistent rotations and chirality-safe flips as the main drivers, with principal-point realignment providing a minor additional benefit.
Beyond monocular detection, we show that the same camera-centric formulation extends to both dense and multi-modal 3D tasks. Plugging 3DRot into the BTS monocular depth estimator yields additional gains on top of its standard augmentation recipe, reducing NYU Depth v2 abs-rel from 0.1783 to 0.1685 and improving $\delta<1.25$ from 0.7472 to 0.7548, while also lowering cross-dataset error on SUN RGB-D. On KITTI, applying 3DRot to MVX-Net (LiDAR+RGB) improves moderate 3D AP from about 63.85 to 65.16 and remains compatible with GlobalRotScaleTrans and other standard 3D augmentations. Together, these results position 3DRot as a simple, depth-free primitive that can be dropped into diverse RGB-based 3D pipelines with minimal changes.
2 Related Work
@@
2.1 RGB-based Rotation Augmentation
@@
Early attempts to regularize RGB-based networks with viewpoint changes started from settings in which the depth channel is readily available, so the augmentation can be expressed directly in camera space. Keypoint-based Monocular 3D Object Detector (Kim et al., 2020) injects small perturbations into the robot‐mounted camera’s roll, pitch and yaw and then re-projects SUN RGB-D boxes to match the new pose. While this improves robustness to head motions, it still relies on full RGB-D frames. Two-level Data Augmentation (Engilberge et al., 2023) pushes consistency a step further: their two-level scheme first applies an in-plane homography to each calibrated view and then corrects the per-view projection matrix so that all images still meet at a common ground plane. The operation is, however, intrinsically coplanar; any 3D structure that rises above or falls below the reference plane (e.g. shelves or hanging objects) accumulates geometric error after transformation. Most recently, GroundMix (Meier et al., 2024) shows that simple in-plane rotations around the optical axis suffice for monocular car- and drone-view detectors; they adjust 3D yaw accordingly and keep the remaining pose angles fixed. This assumption holds only when the camera truly spins about its z-axis; moreover, it offers no rigorous proof of geometric consistency, merely demonstrating visual plausibility.
2.2 RGB-based Flipping Augmentation
@@
Random horizontal flipping is the most widely used geometry-aware transform in RGB-based tasks because it is label‐preserving for the image plane and inexpensive to implement. The pioneering (Brazil and Liu, 2019) pipeline reflects each 3D box corner on the image, then back-projects the flipped vertices with the original depth to obtain an updated 3D cuboid, leaving all other pose parameters. Although effective for heading predictions, this strategy cannot propagate sign changes to the off-diagonal entries of a full $3{\times}3$ rotation matrix and therefore assumes that only the yaw angle is modeled. Several follow-up methods (Chen et al., 2020; Zou et al., 2021; Li et al., 2022, 2024; Wang et al., 2021; Pan et al., 2025a) inherit the same augmentation. Cube R-CNN (Brazil et al., 2023) introduces a more principled update: it composes two mirror matrices, one for the image warp and one for the camera coordinate frame, so that the predicted $R\!\in\!SO(3)$ remains valid after a flip. This design finally enables horizontal mirroring for full‐angle pose regression, but the derivation presumes that the principal point coincides with the image center; when intrinsics are asymmetric or cropped, the 2D-3D alignment still drifts.
2.3 Other Augmentations
@@
Beyond rotations and flips, several geometry–aware strategies have been explored to enrich RGB-based 3D data. Geometric Consistency Augmentation (Lian et al., 2022) introduces four camera-centric manipulations-random scaling, random cropping, camera translation, and instance‐level copy-paste-each designed to preserve depth cues after image warping. While these augmentations improve robustness, random cropping can discard large portions of background context, and the copy-paste pipeline requires dense depth maps plus heuristic rules to blend pasted objects seamlessly. 3D Copy-Paste (Ge et al., 2023) makes this idea more rigorous by sampling physically plausible insertions from an external CAD library: candidate objects are resized, posed, and lit to avoid collisions and match the scene illumination, yielding more realistic composites. Nevertheless, its rendering loop and collision search add notable pre-processing cost, and downstream accuracy still depends on perfect plane reconstruction and light estimation. Moreover, random cropping risks discarding the background context that conveys important geometric cues, while instance-level copy paste-despite its popularity-relies on intricate manual placement heuristics and still falls short of producing truly photorealistic composites (Li et al., 2024).
3 Method
@@
3DRot applies an arbitrary camera‑centric rotation (pitch, roll, yaw; optional mirroring) about the optical center and warps the image using the corresponding pure‑rotation homography. This keeps updates to the camera frame and screen space consistent and preserves the projection geometry. We first outline the framework and notation, then derive the rotation model and label‑synchronization rules. Intuitively, 3DRot rotates each pixel’s viewing ray about the optical center, preserving all geometric correspondences without introducing any new quantities that would need to be re-estimated. A conceptual illustration of this camera-centric rotation and synchronized label update is shown in Fig. 3.
@@
Figure 3: Left: original image; right: view rotated about the optical center with roll $30^{\circ}$. Insets: the ray-imaging diagram (top-right) rotates the image plane about the optical center while keeping the viewing rays fixed, and the pose-imaging diagram (bottom-left) rotates the camera frame so that 3D labels remain valid. The resulting RGB on the right preserves 2D–3D geometric consistency while only the screen footprint changes. See Supplementary Video for an animated demo.
3.1 Coordinate System and 3D Cuboid Transformation
3DRot is implemented as a rigid rotation of the camera coordinate frame about the optical center, followed by re-expressing all scene geometry in the rotated camera frame. In the standard world-to-camera formulation, this amounts to composing the original camera extrinsics with the sampled rotation. For individual 3D points, their coordinates in the camera frame are simply left-multiplied by the sampled camera rotation. For axis-aligned cuboids parameterized by a rotation, side lengths, and a 3D center, the same camera rotation is applied to the cuboid rotation and center, while leaving the side lengths unchanged. Appendix B provides the full derivation starting from the world-to-camera transform and shows how this point-wise update lifts to the cuboid representation used throughout this paper.
3.2 Projective Transformation
Having established how the 3D coordinates of each scene point are updated under a camera-centric rotation, we now turn to their image-plane projections. Specifically, we must determine how every pixel location transforms so that the rotated view remains geometry-consistent with the new camera frame. To this end, we (i) recap the pinhole-projection equation, (ii) derive the homography that arises when the camera rotates about its optical center, and (iii) show how this homography simplifies to a pure-rotation form that is valid for any 3D scene, without the usual planarity assumption or depth information. These steps yield a closed-form mapping $H=K^{\prime}R_{c}K^{-1}$ which we will use to warp the RGB image and masks in the remainder of the pipeline.